# Are We Overlooking the Dimensions? Learning Latent Hierarchical Channel Structure for High-Dimensional Time Series Forecasting

### v1

@@ Appendix B Future Opportunity
Appendix C Limitations
First, U-Cast still needs attention to capture channel correlation, which incurs high computational cost although it has already achieved superior efficiency compared to other baselines. Second, the number of layers and the reduction ratio need to be predefined, which may limit its utilization across datasets with varying hierarchical complexity and scale. Third, the effectiveness of U-Cast relies on the presence of strong latent hierarchical structures in the data. In datasets where inter-channel dependencies are weak, sparse, or not semantically organized, the performance advantage may diminish. Furthermore, the model assumes that these structures are learnable through attention and full-rank regularization, but this may not hold in scenarios with extremely noisy or low-signal input features, leading to suboptimal disentanglement and degraded forecasting accuracy.
Appendix D Social Impact
Transparency Impact. Transparency is essential for promoting accountability, enabling fair comparison, and advancing collective understanding. To support these goals, we will publicly release our codebase, datasets, and evaluation pipeline. However, in accordance with anonymized submission requirements, these materials are currently provided only in the supplementary material.
Environmental Impact. U-Cast offers efficient inference, which reduces GPU runtime and, consequently, the overall energy consumption. This efficiency helps lower the associated carbon emissions, particularly from energy-intensive local power grids.
Ethical Impact. Although U-Cast shows strong performance in high-dimensional time series forecasting, its deployment in critical domains such as healthcare requires caution. In high-stakes settings, inaccurate forecasts could lead to harmful outcomes, underscoring the importance of thorough validation and domain-specific safeguards.
Appendix E Theorem 1 Proof
@@
Again, using the independence of $\varepsilon$ from $w$ and $x$ (and thus from $\mathbb{E}[w\mid x]$), the cross-term expectation is zero: $\mathbb{E}[(w-\mathbb{E}[w\mid x])\varepsilon]=\mathbb{E}[w-\mathbb{E}[w\mid x]]\mathbb{E}[\varepsilon]=0$. The first term is $a_{12}^{2}$ times the definition of the conditional variance, $\operatorname{Var}(w\mid x)=\mathbb{E}[(w-\mathbb{E}[w\mid x])^{2}]$, since $w$ has zero mean (due to normalization) and $\mathbb{E}[w\mid x]$ is the conditional mean. Therefore,
$R_{\text{CI}}$ $=a_{12}^{2}\operatorname{Var}(w\mid x)+\mathbb{E}[\varepsilon^{2}]$
$=\sigma_{11}+a_{12}^{2}\operatorname{Var}(w\mid x).$
@@
$R_{\text{CI}}-R_{\text{CD}}=(\sigma_{11}+a_{12}^{2}\operatorname{Var}(w\mid x))-\sigma_{11}=a_{12}^{2}\operatorname{Var}(w\mid x).$
This difference is non-negative since $a_{12}^{2}\geq 0$ and the conditional variance $\operatorname{Var}(w\mid x)\geq 0$. The risk reduction is strictly positive if $a_{12}\neq 0$ and $\operatorname{Var}(w\mid x)>0$. This occurs when the second channel ($w$) has a non-zero direct influence on the first channel in the next period ($a_{12}\neq 0$) and when the second channel carries some information not already present in the first channel (i.e., $w$ is not perfectly predictable from $x$, ensuring $\operatorname{Var}(w\mid x)>0$). If $a_{12}=0$, or if $w$ is perfectly predictable from $x$, the risks are equal. Similarly, if we focus on forecasting $z_{t+1}^{(2)}$, we have:
$R_{\text{CI}}-R_{\text{CD}}=(\sigma_{22}+a_{21}^{2}\operatorname{Var}(w\mid x))-\sigma_{22}=a_{21}^{2}\operatorname{Var}(x\mid w).$
Thus, $\sum_{i}^{2}(R_{\text{CI}}-R_{\text{CD}})=a_{12}^{2}\operatorname{Var}(w\mid x)+a_{21}^{2}\operatorname{Var}(x\mid w)>0$ This example demonstrates that, under squared-error loss for a linear Gaussian VAR model, using the CD leads to a Bayes risk that is always less than or equal to the Bayes risk obtained using only the CI.
Appendix F Theorem 2 Proof
@@
  • $\operatorname{Var}(z_{t}^{(p)}\mid z_{t}^{(1)},\dots,z_{t}^{(p-1)})>0$, so that $z_{t}^{(p)}$ is not a deterministic function of the previous channels.
@@
Appendix G Theorem 3 Proof
@@
Write the compact singular–value decomposition of $\mathbf{H}^{(\ell-1)}$ as $\mathbf{H}^{(\ell-1)}=\mathbf{U}\,\mathbf{\Sigma}\,\mathbf{V}^{\top},$ where $\mathbf{U}\in\mathbb{R}^{C\times r},\;\mathbf{\Sigma}\in\mathbb{R}^{r\times r}\;\text{is diagonal with positive entries},\;\mathbf{V}\in\mathbb{R}^{d\times r},$ and the columns of $\mathbf{U}$ form an orthonormal basis of the row-space of $\mathbf{H}^{(\ell-1)}$. Because $r\geq C^{\prime}$ by assumption, we can pick a matrix $\mathbf{Q}\in\mathbb{R}^{C^{\prime}\times C}$ of full row rank whose rows are any $C^{\prime}$ independent linear combinations of the rows of $\mathbf{U}^{\top}$. Formally, let $\mathbf{S}\in\mathbb{R}^{C^{\prime}\times C}$ select those $C^{\prime}$ rows (so $\mathbf{S}\mathbf{S}^{\top}=\mathbf{I}_{C^{\prime}}$), and set $\mathbf{Q}=\mathbf{S}$. Then
@@
which proves the first claim. Assume further that $r\geq d$ and $C^{\prime}\geq d$, so the maximal attainable row rank is $d$. Define $\mathbf{\Sigma}^{(\ell)}=\tfrac{1}{d}\mathbf{H}^{(\ell)}\mathbf{H}^{(\ell)\!\top}\in\mathbb{R}^{C^{\prime}\times C^{\prime}}.$ Adding the penalty $\mathcal{L}_{\mathrm{cov}}=-\tfrac{1}{C^{\prime}}\log\det\!\bigl{(}\mathbf{\Sigma}^{(\ell)}+\varepsilon\mathbf{I}_{C^{\prime}}\bigr{)}$ in training maximises $\det\!\bigl{(}\mathbf{\Sigma}^{(\ell)}\bigr{)}=\prod_{i=1}^{C^{\prime}}\sigma_{i}^{2},$ where $\sigma_{i}$ are the singular values of $\mathbf{H}^{(\ell)}$. Because the product of the singular values grows when any $\sigma_{i}$ that is near zero increases, gradient descent on $-\log\det(\cdot)$ pushes all $\sigma_{i}$ away from zero, thereby driving $\mathbf{H}^{(\ell)}$ toward full row rank $\min(C^{\prime},d)$.
@@
Appendix H Theorem of Information Reduction
@@
[…]
@@
˙$\dot{\mathcal{L}}_{\mathrm{cov}}=-\frac{1}{C^{\prime}}\operatorname{tr}\!\Bigl{(}(\mathbf{\Sigma}+\varepsilon I_{C^{\prime}})^{-1}\dot{\mathbf{\Sigma}}\Bigr{)}.$
@@
If the optimisation rule ensures ˙$\dot{\mathcal{L}}_{\mathrm{cov}}\leq 0$ (e.g. gradient descent or any monotone line search), then the inequality gives $\dot{h}\geq 0.$ Equality holds only when ˙$\dot{\mathcal{L}}_{\mathrm{cov}}=0$, which, under differentiability, implies the gradient of the loss is zero. Therefore entropy increases strictly whenever the loss decreases strictly. ∎
Appendix I Detailed Experimental Setup for Empirical Analysis
@@
I.1 Synthetic Data Generation
@@
I.2 Forecasting Task and Data Preparation
@@ I.5 Evaluation Metrics
Appendix J Computational Complexity
@@
$C$ – number of channels (tokens in the channel dimension); $T$ – look-back length (time steps); $d$ – model width; $h$ – number of heads; $r>1$ – reduction ratio; $L=\lceil\log_{r}C\rceil$ – number of encoder/decoder stages. Before the first attention layer, both UCast and iTransformer apply a single linear map $\mathbf{W}_{\text{in}}\!\in\!\mathbb{R}^{T\times d}$ that compresses each length-$T$ series to a $d$-dimensional embedding. Consequently, attention thereafter depends only on channel counts, not on $T$.
@@
At stage $\ell$ the query length is $M_{\ell}=C/r^{\,\ell}$ and the key/value length is $C_{\ell-1}=C/r^{\,(\ell-1)}$. Scaled-dot-product attention therefore costs
@@
$\boxed{\mathcal{T}_{\text{UCast}}=\mathcal{O}\!\Bigl{(}\tfrac{C^{2}d}{h\,r}\Bigr{)}}.$
@@
$\boxed{\mathcal{M}_{\text{UCast}}=\mathcal{O}\!\Bigl{(}\tfrac{C^{2}}{r}\Bigr{)}}.$
@@
iTransformer also compresses the time axis to width $d$ and then applies full self-attention once over the $C$ channel tokens:
$\boxed{\mathcal{T}_{\text{iTrans}}=\mathcal{O}\!\Bigl{(}\tfrac{C^{2}d}{h}\Bigr{)}},\qquad\boxed{\mathcal{M}_{\text{iTrans}}=\mathcal{O}\!\bigl{(}C^{2}\bigr{)}}.$

### v2

@@ Appendix B Future Opportunity
Appendix C Broader Impact
This work reframes time series forecasting (TSF) around high-dimensional TSF (HDTSF), where models must reason over thousands of channels and their structure. By curating Time-HD across more than ten domains and large channel counts, we move beyond the past practice of training and testing only on small, low-dimensional benchmarks. This shift makes evaluation more reliable, reveals failure modes that only appear at scale, and brings assessment closer to real deployments in sensor networks, markets, and geoscience. It also encourages the community to measure real gains rather than benchmark-specific gains.
Time-HD and Time-HD-Lib are released to lower the barrier to entry and raise reproducibility. The dataset suite standardizes timestamps, horizons, and metrics; the library offers a distributed train–validate–test pipeline with common preprocessing and reference configs. Together, they let researchers and practitioners start quickly, scale experiments without bespoke infrastructure, and add new datasets and tasks that broaden HDTSF coverage. This shared stack supports fair comparisons and faster iteration on modeling ideas.
Methodologically, U-Cast serves as a strong baseline for future work on HDTSF. Its design targets cross-channel structure and efficiency, yielding competitive accuracy with favorable speed and memory use. We recommend using U-Cast as a starting point for ablations and new architectures, reporting uncertainty alongside point forecasts, and disaggregating results by domain. Users should watch for dataset bias and compute cost; the released pipeline helps by making runs reproducible and efficient. Overall, the task, resources, and baseline aim to help the TSF community build and evaluate models that scale to realistic settings.
Appendix D Datasets Description
Temp (Global Temp) and Wind (Global Wind). We adapt this dataset from Corrformer (Wu et al., 2023). Global Temp&Wind dataset is from the National Centers for Environmental Information (NCEI)  [^2] https://www.ncei.noaa.gov/. This dataset contains the hourly averaged wind speed and hourly temperature of 3,850 stations (7,700 dimensions) around the world from 01/01/2019 to 12/31/2020, 17,544 timesteps in total. This dataset is well established. However, we find there are some constant columns in the datasets, which means the temp or wind speed never change. We remove these columns.
Solar (NREL Solar Power). NREL Solar Power is collected from National Renewable Energy Laboratory (NREL)  [^3] https://www.nrel.gov/grid/solar-power-data.html. One of the most popular datasets for time series forecasting is Solar-Energy (Lai et al., 2018), which is sourced from NREL. However, it includes data from only 137 solar photovoltaic (PV) plants in Alabama, limiting the dataset to 137 dimensions. To increase the dimensionality, we collect data from 5,166 PV plants across 47 U.S. states, excluding Alaska, Hawaii, and North Dakota, using data from NREL. The data source is clean so we don’t conduct further preprocessing on it.
Traffic-CA (LargeST-CA), Traffic-GLA (LargeST-GLA), Traffic-GBA (LargeST-GBA). The source of this dataset is the Caltrans Performance Measurement System (PeMS) [^4] https://pems.dot.ca.gov/, and we utilize its preprocessed version(Liu et al., 2023). LargeST-CA encompassing a total of 8,600 sensors in California and each sensor contains five years of traffic flow data (from 2017 to 2021) with a 5-minute interval, resulting in a total of 525,888 time frames. LargeST-CA consists of three subsets: (1) LargeST-GLA includes 3,834 sensors installed across five counties in the Greater Los Angeles (GLA) area: Los Angeles, Orange, Riverside, San Bernardino, and Ventura. (2) LargeST-GBA contains 2,352 sensors distributed across 11 counties in the Greater Bay Area (GBA): Alameda, Contra Costa, Marin, Napa, San Benito, San Francisco, San Mateo, Santa Clara, Santa Cruz, Solano, and Sonoma. (3) LargeST-SD, the smallest subset, consists of 716 sensors located exclusively in San Diego County. Since LargeST-SD has only 716 dimensions, we do not include it in our benchmark. For the 525,888 original time frames, we aggregate the data into 1-hour intervals, resulting in 43,824 time frames. We then remove sensors with more than 100 missing (NaN) values, and apply linear interpolation to fill the remaining missing values.
Neurolib. Simulations made use of the Neurolib code (Cakan et al., 2021). We utilize the brain structure in ABIDE dataset (Craddock et al., 2013).
SIRS. Simulations extend the classical SIR model (Cai et al., 2015) by incorporating temporary immunity, resulting in dynamic transitions among Susceptible, Infected, and Recovered groups within a single region. To enhance realism, parameter noise, and seasonal variation are included. For multiple regions, each compartment has a fixed transfer rate between regions to simulate commuting and spatial transmission dynamics.
Air Quality (Chinese Air Quality). This dataset includes 7 air quality indicators: $AQI$, COCO, $NO_{2}$, $O_{3}$, $PM10$, $PM2.5$, and $SO_{2}$ of 786 distinct cities in China. We remove dimensions with more than 100 missing values and interpolate other dimensions linearly, resulting in 1,106 dimensions.
M5. We adapt this dataset from a Kaggle competition (Howard et al., 2020). It covers stores in three U.S. states, i.e. California, Texas, and Wisconsin, and includes item-level sales data, department and product category details, and store information. The dataset records daily sales for each item from 07/01/2015 to 06/30/2022 (a total of 1,947 days) across different departments and stores, resulting in 30,490 time series. Due to its sparsity (i.e., a large number of zero values), we aggregate sales by summing each item’s sales across all departments and stores, reducing the dataset to 3,049 aggregated items.
Measles (Measles England). The measles dataset (Madden et al., 2024; Liu et al., 2025; 2024b), contains biweekly measles infections from 387 regions across England and Wales from 1944–1965. For each region, the dataset includes three features: inferred infections, population, and suspected cases, resulting in 1,161 dimensions.
Wiki-20k (Wikipedia Web Traffic). The dataset is adapted from the Monash Time Series Forecasting Repository  [^8] https://zenodo.org/records/7370977. It comprises 145,063 time series representing web traffic for various Wikipedia pages from July 1, 2015, to June 30, 2022 (2,557 days). This dataset is an extended version of the one used in the Kaggle Wikipedia Web Traffic Forecasting Competition  [^9] https://www.kaggle.com/c/web-traffic-time-series-forecasting/data. Due to data sparsity, we remove the Wikipedia pages containing NaN values, resulting in a final dataset of 112,333 Wikipedia pages, each corresponding to one dimension. Due to the original dataset’s dimension is too large and is not correlated, we select the first 20k dimensions to construct Wiki-20k dataset for evaluation.
Mobility (Google Community Mobility). The Community Mobility dataset (Google, 2020) aims to assist in analyzing the impact of COVID-19 on community movement patterns. The dataset covers movement trends over time (from February 15, 2020, to October 15, 2022, spanning 974 days) across 4,334 regions in different countries. It categorizes mobility data into six place types: retail and recreation (restaurants, cafes, shopping centers, theme parks, museums, libraries, and movie theaters), groceries and pharmacies (grocery stores, food warehouses, farmers markets, specialty food shops, drug stores, and pharmacies), parks (local and national parks, public beaches, marinas, dog parks, plazas, and public gardens), transit stations (subway, bus, and train stations), workplaces (places of employment), and residential (places of residence). The dataset initially consists of 26,004 dimensions, derived from the six place categories across 4,334 regions. However, due to data sparsity, we exclude regions containing NaN values, reducing the dataset to 971 regions and 5,826 dimensions. The values in the dataset represent changes in visit frequency and duration of stay at various locations relative to the baseline (i.e. the median value for the same day of the week during the five-week period from January 3 to February 6, 2020.). These changes are computed using aggregated and anonymized data, similar to the methodology employed in Google Maps to estimate popular visit times.
Prediction Length $S$
Earlier benchmarks commonly evaluated the same four horizons (96, 192, 336, 720) across all datasets, regardless of sampling frequency. This practice can yield scenarios with limited practical relevance, for example, requiring a weather model to predict 720 days ahead. In contrast, we align the prediction horizon with the data resolution so that each task corresponds to a realistic operational need:
  • 1 ms–30 min data $\to\;$ $S=336$ steps (0.34 s to 28 h)
  • 1 h data $\to\;$ $S=168$ steps (7 days)
  • 6 h data $\to\;$ $S=28$ steps (7 days)
  • 1 day data $\to\;$ $S=7$ steps (1 week)
These settings, listed under “Pred Length” in Table 3, reflect forecasting horizons commonly adopted by domain experts (e.g., weekly planning in retail or epidemiology).
Appendix E Analysis of Channel Correlation in Time-HD
We follow TFB (Qiu et al., 2024a) to compute channel correlation. In multivariate time series, different variables often share common temporal patterns due to underlying shared factors. To quantify this, we extract 22 features per channel using Catch22 (Lubba et al., 2019). These features serve as fixed-length representations of each time series channel. We then compute the pairwise Pearson correlation coefficients (Cohen et al., 2009) between all channel representations and report the mean and variance.
The results reported in the “Correlation” column of Table 3 show that all datasets exhibit high average channel correlation, reflecting strong inter-variable dependencies and a dominant global component shared across channels. While this global component is highly predictable, it provides little additional information for forecasting, since it largely repeats patterns common to all variables. When such shared trends dominate, they can overshadow the subtle yet critical channel-specific dynamics that ultimately determine predictive accuracy. Consequently, a model that focuses primarily on global trends will miss the fine-grained deviations that differentiate one channel from another. These findings motivate the need for models that explicitly disentangle and separate the global component from channel-specific signals while also capturing structured inter-channel dependencies.
Modeling Strategy in U-Cast.
U-Cast addresses this challenge through an architecture that separates the modeling of unique channel dynamics from that of shared global trends (see Figure 1). Specifically:
  • Modeling Hierarchical Inter-Channel Dependency: The Hierarchical Latent Query Network compresses the input into a low-dimensional latent space, where the Full-Rank Regularization (Section 4.2) enforces disentanglement and removes redundancy. The subsequent Temporal Alignment module then captures the essential channel-specific temporal patterns within this compact space.
  • Modeling Shared Trends: After unique inter-channel dynamics are learned, the Hierarchical Upsampling Network and final Output Projection layer (Equation 6) reconstruct the forecasts for all channels. With a residual connection to the original embeddings, the projection layer re-applies the simple global trends on top of the channel-specific predictions from the latent space.
In essence, U-Cast decomposes the problem: it compresses and disentangles redundant global components to focus on informative inter-channel dynamics, while leveraging the output projection to efficiently recover global trends. This separation is not a loss of information but a deliberate design choice that enables accurate and robust forecasting in high-dimensional, highly correlated time series.
We also want to highlight that, in the real world, there indeed exist time series collections with low inter-channel correlations. However, merging such weakly related time series into a single dataset is usually not meaningful because they do not provide complementary predictive information to each other. In contrast, the datasets in Time-HD are specifically collected from scenarios with strong inter-channel correlations, where exploiting these correlations can meaningfully improve forecasting performance.
Appendix F Implementation Details
Metric Details. We use Mean Square Error (MSE) and Mean Absolute Error (MAE) as our evaluation metrics, following (Liu et al., 2024a; Luo & Wang, 2024; Wang et al., 2024a; Wu et al., 2022; 2021; Zhou et al., 2021; Zeng et al., 2023):
Experiment Details. All experiments are implemented in PyTorch (Paszke, 2019) and conducted on a cluster equipped with 8 NVIDIA A100 GPUs and several NVIDIA H100 GPUs. The baselines are trained using their default configurations as reported in their respective papers, and with further hyperparameter searching (Learning Rate in {0.01, 0.001, 0.0001}, Input Sequence Length $T$ in $3\times S$, $4\times S$, $5\times S$}), where $S$ is prediction length. ADAM optimizer (Kingma & Ba, 2014) is used with MSE loss for the training of all the models. We apply early stopping with a patience value of 5 epochs. The batch size is initially set to 32. If an out-of-memory (OOM) error occurs, the batch size is automatically halved until the issue is resolved. For U-Cast, we default to using non-stationary (Liu et al., 2022) normalization for all datasets. We default to set the number of layers $L$ to 2, the hidden dimension $d$ to 512, and the reduction ratio $r$ to 16 for all the datasets. Additional detailed U-Cast model configuration information is presented in Table 8. We conduct hyperparameter sensitivity analysis in Appendix O. For robustness, we perform five-seed runs on all datasets and calculate the standard deviation. As shown in Table 7, U-Cast trains reliably across all cases.
Table 7: Performance of U-Cast (five-seed runs) on all datasets. Results are reported as mean $\pm$ standard deviation.
Dataset MSE MAE
Atec $0.290\pm 0.001$ $0.278\pm 0.002$
Air Quality $0.420\pm 0.005$ $0.411\pm 0.004$
Temp $0.263\pm 0.002$ $0.384\pm 0.001$
Wind $1.078\pm 0.003$ $0.686\pm 0.002$
Mobility $0.317\pm 0.004$ $0.324\pm 0.003$
Traffic-CA $0.060\pm 0.001$ $0.134\pm 0.002$
Traffic-GBA $0.059\pm 0.000$ $0.128\pm 0.000$
Traffic-GLA $0.059\pm 0.000$ $0.135\pm 0.001$
M5 $3.472\pm 0.001$ $0.843\pm 0.000$
Measles $014\pm 0.005 0.0$$050\pm 0.004$
Neurolib $1.747\pm 0.005$ $0.349\pm 0.006$
Solar $0.159\pm 0.001$ $0.220\pm 0.002$
SIRS $006\pm 0.000 0.0$$048\pm 0.001$
Meters $0.947\pm 0.011$ $0.550\pm 0.009$
SP500 $0.501\pm 0.008$ $0.305\pm 0.006$
Wiki-20k $10.275\pm 0.005$ $0.302\pm 0.004$
Table 8: U-Cast hyperparameters used for each dataset.
Dataset Learning Rate Input Sequence Length $T$ $\alpha$
Atec 0.0005 $3\times S$ 0.001
Air Quality 0.0005 $3\times S$ 0.100
Temp 0.001 $3\times S$ 0.001
Wind 0.0005 $4\times S$ 0.010
Mobility 0.0005 $4\times S$ 0.010
Traffic-CA 0.002 $4\times S$ 0.010
Traffic-GBA 0.002 $4\times S$ 0.010
Traffic-GLA 0.002 $4\times S$ 0.010
M5 0.001 $4\times S$ 10.000
Measles 0.0005 $4\times S$ 0.010
Neurolib 0.0005 $4\times S$ 0.010
Solar 0.001 $4\times S$ 0.010
SIRS 0.001 $4\times S$ 0.001
Meters 0.0005 $4\times S$ 0.010
SP500 0.002 $4\times S$ 0.001
Wiki-20k 0.001 $4\times S$ 0.001
Appendix G Theorem 1 Proof
@@
Again, using the independence of $\varepsilon$ from $w$ and $x$ (and thus from $\mathbb{E}[w\mid x]$), the cross-term expectation is zero: $\mathbb{E}[(w-\mathbb{E}[w\mid x])\varepsilon]=\mathbb{E}[w-\mathbb{E}[w\mid x]]\mathbb{E}[\varepsilon]=0$. The first term is $a_{12}^{2}$ times the definition of the conditional variance, $\mathrm{Var}(w\mid x)=\mathbb{E}[(w-\mathbb{E}[w\mid x])^{2}]$, since $w$ has zero mean (due to normalization) and $\mathbb{E}[w\mid x]$ is the conditional mean. Therefore,
$R_{\text{CI}}$ $=a_{12}^{2}\mathrm{Var}(w\mid x)+\mathbb{E}[\varepsilon^{2}]$
$=\sigma_{11}+a_{12}^{2}\mathrm{Var}(w\mid x).$
@@
$R_{\text{CI}}-R_{\text{CD}}=(\sigma_{11}+a_{12}^{2}\mathrm{Var}(w\mid x))-\sigma_{11}=a_{12}^{2}\mathrm{Var}(w\mid x).$
This difference is non-negative since $a_{12}^{2}\geq 0$ and the conditional variance $\mathrm{Var}(w\mid x)\geq 0$. The risk reduction is strictly positive if $a_{12}\neq 0$ and $\mathrm{Var}(w\mid x)>0$. This occurs when the second channel ($w$) has a non-zero direct influence on the first channel in the next period ($a_{12}\neq 0$) and when the second channel carries some information not already present in the first channel (i.e., $w$ is not perfectly predictable from $x$, ensuring $\mathrm{Var}(w\mid x)>0$). If $a_{12}=0$, or if $w$ is perfectly predictable from $x$, the risks are equal. Similarly, if we focus on forecasting $z_{t+1}^{(2)}$, we have:
$R_{\text{CI}}-R_{\text{CD}}=(\sigma_{22}+a_{21}^{2}\mathrm{Var}(w\mid x))-\sigma_{22}=a_{21}^{2}\mathrm{Var}(x\mid w).$
Thus, $\sum_{i}^{2}(R_{\text{CI}}-R_{\text{CD}})=a_{12}^{2}\mathrm{Var}(w\mid x)+a_{21}^{2}\mathrm{Var}(x\mid w)>0$ This example demonstrates that, under squared-error loss for a linear Gaussian VAR model, using the CD leads to a Bayes risk that is always less than or equal to the Bayes risk obtained using only the CI.
Appendix H Theorem 2 Proof
@@
  • $\mathrm{Var}(z_{t}^{(p)}\mid z_{t}^{(1)},\dots,z_{t}^{(p-1)})>0$, so that $z_{t}^{(p)}$ is not a deterministic function of the previous channels.
@@
Appendix I Theorem 3 Proof
@@
Write the compact singular–value decomposition of $\mathbf{H}^{(\ell-1)}$ as $\mathbf{H}^{(\ell-1)}=\mathbf{U}\,\mathbf{\Sigma}\,\mathbf{V}^{\top},$ where $\mathbf{U}\in\mathbb{R}^{C\times r},\;\mathbf{\Sigma}\in\mathbb{R}^{r\times r}\;\text{is diagonal with positive entries},\;\mathbf{V}\in\mathbb{R}^{d\times r},$ and the columns of $\mathbf{U}$ form an orthonormal basis of the row‑space of $\mathbf{H}^{(\ell-1)}$. Because $r\geq C^{\prime}$ by assumption, we can pick a matrix $\mathbf{Q}\in\mathbb{R}^{C^{\prime}\times C}$ of full row rank whose rows are any $C^{\prime}$ independent linear combinations of the rows of $\mathbf{U}^{\top}$. Formally, let $\mathbf{S}\in\mathbb{R}^{C^{\prime}\times C}$ select those $C^{\prime}$ rows (so $\mathbf{S}\mathbf{S}^{\top}=\mathbf{I}_{C^{\prime}}$), and set $\mathbf{Q}=\mathbf{S}$. Then
@@
which proves the first claim. Assume further that $r\geq d$ and $C^{\prime}\geq d$, so the maximal attainable row rank is $d$. Define $\mathbf{\Sigma}^{(\ell)}=\tfrac{1}{d}\mathbf{H}^{(\ell)}\mathbf{H}^{(\ell)\!\top}\in\mathbb{R}^{C^{\prime}\times C^{\prime}}.$ Adding the penalty $\mathcal{L}_{\mathrm{cov}}=-\tfrac{1}{C^{\prime}}\log\det\!\bigl(\mathbf{\Sigma}^{(\ell)}+\varepsilon\mathbf{I}_{C^{\prime}}\bigr)$ in training maximises $\det\!\bigl(\mathbf{\Sigma}^{(\ell)}\bigr)=\prod_{i=1}^{C^{\prime}}\sigma_{i}^{2},$ where $\sigma_{i}$ are the singular values of $\mathbf{H}^{(\ell)}$. Because the product of the singular values grows when any $\sigma_{i}$ that is near zero increases, gradient descent on $\log\det(\cdot)$ pushes all $\sigma_{i}$ away from zero, thereby driving $\mathbf{H}^{(\ell)}$ toward full row rank $\min(C^{\prime},d)$.
@@
Appendix J Theorem of Information Reduction
@@
$\frac{\mathrm{d}}{\mathrm{d}t}h^{(\ell)}(t)=-\frac{C^{\prime}}{2}\,\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{L}_{\mathrm{cov}}^{(\ell)}(t)\;\geq\;0,$
@@
$\dot{\mathcal{L}}_{\mathrm{cov}}=-\frac{1}{C^{\prime}}\operatorname{tr}\!\Bigl((\mathbf{\Sigma}+\varepsilon I_{C^{\prime}})^{-1}\dot{\mathbf{\Sigma}}\Bigr).$
@@
If the optimisation rule ensures $\dot{\mathcal{L}}_{\mathrm{cov}}\leq 0$ (e.g. gradient descent or any monotone line search), then the inequality gives $\dot{h}\geq 0.$ Equality holds only when $\dot{\mathcal{L}}_{\mathrm{cov}}=0$, which, under differentiability, implies the gradient of the loss is zero. Therefore entropy increases strictly whenever the loss decreases strictly. ∎
Appendix K Detailed Experimental Setup for Empirical Analysis
@@
K.1 Synthetic Data Generation
@@
K.2 Forecasting Task and Data Preparation
@@
Appendix L Computational Complexity
@@
$C$ – number of channels (tokens in the channel dimension); $T$ – look‑back length (time steps); $d$ – model width; $h$ – number of heads; $r>1$ – reduction ratio; $L=\lceil\log_{r}C\rceil$ – number of encoder/decoder stages. Before the first attention layer, both UCast and iTransformer apply a single linear map $\mathbf{W}_{\text{in}}\!\in\!\mathbb{R}^{T\times d}$ that compresses each length‑$T$ series to a dd‑dimensional embedding. Consequently, attention thereafter depends only on channel counts, not on $T$.
@@
At stage $\ell$ the query length is $M_{\ell}=C/r^{\,\ell}$ and the key/value length is $C_{\ell-1}=C/r^{\,(\ell-1)}$. Scaled‑dot‑product attention therefore costs
@@
$\mathcal{T}_{\text{UCast}}=\mathcal{O}\!\Bigl(\tfrac{C^{2}d}{h\,r}\Bigr).$
@@
$\mathcal{M}_{\text{UCast}}=\mathcal{O}\!\Bigl(\tfrac{C^{2}}{r}\Bigr).$
@@
iTransformer also compresses the time axis to width $d$ and then applies full self‑attention once over the $C$ channel tokens:
$\mathcal{T}_{\text{iTrans}}=\mathcal{O}\!\Bigl(\tfrac{C^{2}d}{h}\Bigr),\qquad\mathcal{M}_{\text{iTrans}}=\mathcal{O}\!\bigl(C^{2}\bigr).$